# Faraday laser pumped cesium beam clock

### v1

@@ Abstract
We realize a high-performance compact optically pumped cesium beam clock using Faraday laser simultaneously as pumping and detection lasers. The Faraday laser, which is frequency stabilized by modulation transfer spectroscopy (MTS) technique, has narrow linewidth and superior frequency stability. Measured by optical heterodyne method between two identical systems, the linewidth of the Faraday laser is 2.5 kHz after MTS locking, and the fractional frequency stability of the Faraday laser is optimized to . Based on this high-performance Faraday laser, the cesium beam clock realizes a signal-to-noise ratio (SNR) in 1 Hz bandwidth of when the cesium oven temperature is 130°C. Frequency-compared with Hydrogen maser, the fractional frequency stability of the Faraday laser pumped cesium beam clock can reach and drops to at 10000 s when the cesium oven temperature is 110°C. This Faraday laser pumped cesium beam clock demonstrates its excellent performance, and its great potential in the fields of timekeeping, navigation, and communication. Meanwhile, the Faraday laser, as a high-performance optical frequency standard, can also contribute to the development of other applications in quantum metrology, precision measurement and atomic physics.
@@ I Introduction
Since being constructed[1], cesium beam clocks have been widely used in the field of frequency and time, promoting the development of timekeeping, high-speed communication, navigation and fundamental research[2, 3]. There are two main types of cesium beam clocks, magnetic state selecting cesium beam clock[4, 5, 6, 7] and optically pumped cesium beam clock[8, 9, 10, 11, 12, 13, 14, 15, 16]. Compared with magnetic state selecting cesium beam clocks, optically pumped cesium beam clocks have a significantly higher atomic utilization, leading to better frequency stability and longer service life. The optically pumped cesium beam clocks is gradually showing a dominant trend.
@@
The frequency stability of optically pumped cesium beam clocks can be further optimized mainly by reducing the shot noise of cesium atoms and the frequency noise of the laser. Increasing the number of interacting atoms can effectively reduce the shot noise of cesium atoms, which can be realized by increasing the atomic beam flux. In practice, the atomic beam flux is chosen to an appropriate finite value, taking into account the lifetime of the cesium clock. When the atomic beam flux is constant, the frequency stability of optically pumped cesium atomic clocks can be greatly improved by adopting a laser with narrow linewidth, low frequency noise and high power stability, which can effectively improve the signal-to-noise ratio (SNR) of Ramsey fringes of the optically pumped cesium beam clock[17]. At the same time, miniaturization and transportability also need to be considered. The volume of the cesium beam clock is mainly limited by the cesium beam tube, but the shortening of the free drift region will reduce the Ramsey evolution time, which will widen the linewidth of the Ramsey fringes to a certain extent and leads to the the deterioration in frequency stability of cesium beam clock. Thus, the optimization of SNR can make the cesium beam clock still have excellent performance. Therefore, lasers with narrow linewidth and low frequency noise are essential.
@@
In order to optimize the frequency stability of optically pumped cesium beam clocks, we use frequency-stabilized Faraday laser [18, 19, 20, 21, 22, 23] as the pumping and detection laser sources. Compared with traditional external-cavity semiconductor lasers (ECDLs), such as lasers frequency selected by interferences and gratings, Faraday laser has the advantage of immunity to temperature and current variations of the laser diode, and the frequency of Faraday laser can be automatically corresponding to atomic transition lines as soon as turning on, since Faraday laser is frequency selected by atomic filter[18, 19, 20]. Moreover, the modulation transfer spectroscopy (MTS) technique[24, 25, 26] is applied to lock the Faraday laser. Compared with other frequency stabilization methods, such as the saturation absorption spectroscopy (SAS)[27, 28, 29], polarization spectroscopy (PS)[30, 31], dichroic atomic vapor laser lock (DAVLL)[32, 33], MTS technique effectively suppress low-frequency noise, and it has the characteristics of high SNR and background-free[24].
@@
In this work, we achieve a high-performance Faraday laser pumped cesium beam clock, which applies an 852 nm Faraday laser frequency-stabilized by MTS technique as pumping and detection lasers. We measure the performance of the frequency-stabilized Faraday laser using the optical heterodyne method, and the result shows that the fractional frequency stability of each system is at 1 s, The most probable linewidth of the heterodyne beating signal is 12.7 kHz when Faraday lasers are free-running, which means that of the single system is 9.0 kHz. The most probable linewidth is 3.5 kHz with Faraday lasers being frequency locked by MTS, thus that of the single system is 2.5 kHz. We use such Faraday laser with narrow linewidth and low frequency noise as the pumping and the detection lasers to build the compact optically pumped cesium beam atomic clock, leading to SNR of the Ramsey fringe being optimized. Thus, the fractional frequency stability of the Faraday laser pumped cesium beam clock can achieve at 1 s and at 100 s, dropping to at 10000 s when the cesium oven temperature is 110°C.
@@ II Experiment
Figure 1 shows the overall structure of the experiment. The Faraday laser is an integrated system with MTS technique for frequency stabilization, and all components are integrated onto a single aluminum base. The Faraday laser is beam-split and frequency-shifted to serve as the pumping and detection lasers for the the cesium beam clock.
@@
The structure of the Faraday laser is shown in Fig. 1. We use the anti-reflection coated laser diode (ARLD) as gain medium. The anti-reflection coating superbly suppresses the influence of the internal cavity mode on the laser mode. Two PBSs, a cesium vapor cell, and permanent magnets compose the Faraday anomalous dispersion optical filter (FADOF). The vapor cell is filled with cesium atoms and 10 torr argon buffer gas, with a diameter of 1.5 cm and a length of 3 cm. Under the action of the magnetic field, the energy levels of the cesium atoms in the vapor cell undergo Zeeman splitting, which gives rise to the Faraday effect. The polarization direction of the linearly polarized light passing through the first PBS rotates as it propagates through the cesium vapor cell, and ultimately only the light whose polarization direction has rotated by can pass through the second PBS. Thus, only a very narrow frequency range of light corresponding to the spectral lines of the atoms can pass through the FADOF. The cesium vapor cell is surrounded with a twisted copper heating wire driven by an heater temperature controller and enclosed by thermal insulation materials. The temperature of the cesium vapor cell is controlled at 66°C with a temperature fluctuation of 0.1°C. The permanent magnets are placed around the cesium vapor cell to form a 900 G longitudinal magnetic field. After collimating by a lens, the fluorescence emitted from the laser diode passes through a FADOF for mode selection and then is reflected back to the laser diode by a high reflector to form the external cavity feedback. Due to the use of ARLD and the extremely narrow transmission spectrum of the FADOF, Faraday lasers are immune to the changes in current and temperature, a characteristic that distinguishes them from other types of external-cavity diode lasers.
@@
In MTS system, the frequency of the Faraday laser is locked by MTS technique to the transition of , as shown in Fig. 1. The Faraday laser beam is divided into pump laser and probe laser for MTS locking by a half wave plate (HWP) and a PBS. The pump laser is modulated by an electro-optic modulator (EOM) with a modulation frequency of 4.6 MHz. Then the modulation of the pump laser is transferred to the porbe laser, which is received by the a photoelectric detector (PD). The electrical signal of the PD is first amplified by an amplifier and then demodulated by a mixer with a demodulation frequency of 4.6 MHz to derive the MTS signal as shown in Fig. 2. The MTS error signal is processed by a proportional integral derivative (PID) circuit (Vescent D2-125) to form a feedback signal, which is send to the laser controller and the laser diode respectively for frequency stabilization. The former signal is processed by the laser controller and then fed back to the current and piezoelectric ceramic (PZT) of the Faraday laser.
@@
As shown in Fig. 1, we implement two identical 852 nm frequency-stabilized Faraday laser systems to realize optical heterodyne measurement. When both faraday lasers are free-running, the signal derived by PD is sent into the spectrum analyzer, then we can obtain the fitted full width at half maximum (FWHM) of the heterodyne beating signal, as shown in Fig. 3(a). Several sets of data are measured under the same conditions to get the most probable linewidth of the free-running Faraday laser which is statistically calculated to be 12.7 kHz, so that the most available linewidth for a single laser was 9.0 kHz, as shown in Fig. 3(b). When both laser systems are locked, the linewidth of the 852 nm MTS systems is obtained by the same method. The corresponding heterodyne beating signal obtained is shown in Fig. 3(c). 60 sets of data are measured under the same conditions, and the most probable linewidth is reduced to 3.5 kHz, so the most allowable linewidth of a single 852 nm MTS systems is 2.5 kHz, as shown in Fig. 3(d). The significant narrowing effect of the Faraday laser linewidth is caused by the MTS technique, which greatly reduces the noise of the free-running Faraday laser. The narrow linewidth characteristic of the frequency-stabilized Faraday laser can greatly expand its application range.
@@
At the same time, the heterodyne beating signal from the PD is sent into a frequency counter (Keysight 53230A) to measure the frequency stability of the Farday laser. The frequency stability of the free-running Faraday laser measured by heterodyne method is at 1 s, as shown in Fig. 4. With both laser systems locked, the corresponding caculated frequency stability is shown in Fig. 4. The frequency stability of the heterodyne beating signal is at 1 s, so the frequency stability of the single 852 nm MTS system is at 1 s. In the MTS system, we use a double-layer vapor cell as the absolute frequency reference. Changes in the temperature of the atomic vapor cell bring about a drift in the frequency of the cesium transition lines, whereas the space between the two quartz layers of the double-layer vapor cell is pumped into a vacuum, and the double-layer vapor cell is therefore well insulated from the effect of ambient temperature fluctuations on the frequency of the cesium transition lines.
@@
In the compact optically pumped cesium beam atomic clock, the pumping and the detection lasers are beam-split from the same Faraday laser frequency-stabilized on the cesium cycling transition (). The frequency-stabilized Faraday laser is divided into two beams. One beam is directly used as detection laser and the other beam is frequency-shifted by 251.1 MHz with a acousto-optic modulator (AOM) and then used as the pumping laser which has a frequency corresponding to the transition line. The corresponding atomic transitions is schematically shown in Fig 1. Both the pumping and the detection lasers pass through the beam expanders before being injected into the cesium beam tube to expand the interaction area between the optical beam and the cesium beam by a factor of . In the area where the cesium beam interacts with the pumping laser, a reflector is mounted on the back of the cesium beam tube to reflect the pumping laser to interact with the cesium beam again to improve the pumping efficiency.
@@ III Results
A microwave signal synthesized by a 10 MHz voltage-controlled crystal oscillator (VCXO) is acted on the microwave cavity and tuned with a center frequency of 9.192631770 GHz. The cesium oven temperature is set at 110 °C. The linewidth and the amplitude of the Ramsey fringe is 551.4 Hz and 461.85 mV, as shown in Fig. 5. The cesium beam clock operates at a modulation frequency of 137 Hz, and the noise of the Ramsey fringe signal is analyzed by a fast Fourier transform spectrum analyzer (Stanford Research Systems, SR770). The noise at 137 Hz is . Thus, the SNR in 1 Hz bandwidth at 137 Hz is given by , where represents the amplitude of the Ramsey fringe and represents the noise of the Ramsey fringe at 137 Hz. The Ramsey fringe signal is modulated and demodulated to produce an error signal which is used to stabilize the output frequency of the VCXO by the servo circuits.
After locking the 10 MHz VCXO of the cesium clock, we compare 10 MHz output of the VCXO with a hydrogen maser (VREMYA-CH, VCH-1003M, option L) by a high-performance phase noise and Allan deviation analyzer (Microsemi, 5120A). The frequency stability of the Faraday laser-based cesium beam clock is two times better than that of the DFB laser-based cesium clock[34], as shown in Fig. 6. In addition, the frequency stability of our cesium clock is also significantly better than that of the cesium clock based on the interferometer ECDL[13], which is at 100 s. The three cesium clocks are different only in the categories of the lasers. That is, they use the same cesium beam tube and electronic system, and thier cesium oven temperatures are all set at 100°C. Thus, the comparison results highlight the superiority of the Faraday laser, which has narrow linewidth and low frequency noise.
We try different cesium oven temperatures, which would directly affect the SNR of the Ramsey fringe. As shown in Fig. 7, when the cesium oven temperature is increased to 110°C, the frequency stability of the cesium beam clock reahces , decreasing to at 10000 s. The frequency stability of our clock is better than that of any reported compact cesium beam atomic clocks to our knowledge, which is attributed to the narrow linewidth and the low frequency noise of the Faraday laser system resulting in high SNR of the Ramsey fringe. Meanwhile, the frequency stability of our clock is at 1 s , which firstly reaches the magnitude. When the cesium oven temperature is 130°C, the SNR of the Ramsey fringe in 1 Hz bandwidth at 137 Hz will reach , which can result in better frequency stability. Taking into account the lifetime of the cesium beam tube, the oven temperature is set at 110°C. The increase in signal intensity brought about by the increase in oven temperature does not necessarily lead directly to a higher SNR, which requires the support of a laser with low frequency noise[35].

### v2

@@ Abstract
We realize a high-performance compact optically pumped cesium beam atomic clock utilizing Faraday laser simultaneously as pumping and detection lasers. The Faraday laser, frequency stabilized by modulation transfer spectroscopy (MTS) technique, has narrow linewidth and superior frequency stability. Measured by optical heterodyne method between two identical systems, the linewidth of the Faraday laser is 2.5 kHz after MTS locking, and the fractional frequency stability of the Faraday laser is optimized to . Based on this high-performance Faraday laser, the cesium beam clock realizes a signal-to-noise ratio (SNR) in 1 Hz bandwidth of when the cesium oven temperature is 130°C. Frequency-compared with Hydrogen maser, the fractional frequency stability of the Faraday laser pumped cesium beam clock can reach and decreases to at 10000 s with the cesium oven temperature set at 110°C. This Faraday laser pumped cesium beam clock demonstrates its excellent performance, and its great potential in the fields of timekeeping, navigation, and communication. Meanwhile, the Faraday laser, as a high-performance optical frequency standard, can also contribute to the development of other applications in quantum metrology, precision measurement, and atomic physics.
@@ I Introduction
Since being constructed[1], cesium beam clocks have been widely used in the field of frequency and time, promoting the development of timekeeping, high-speed communication, navigation, and fundamental research[2, 3]. There are two main types of cesium beam clocks, magnetic state selecting cesium beam clock[4, 5, 6, 7] and optically pumped cesium beam clock[8, 9, 10, 11, 12, 13, 14, 15, 16]. Compared with magnetic state selecting cesium beam clocks, optically pumped cesium beam clocks have a significantly higher atomic utilization, leading to better frequency stability and longer service life. The optically pumped cesium beam clocks is gradually showing a dominant trend.
@@
The frequency stability of optically pumped cesium beam clocks can be further optimized mainly by reducing the shot noise of cesium atoms and the frequency noise of the laser. Increasing the number of interacting atoms can effectively reduce the shot noise of cesium atoms, which can be realized by increasing the atomic beam flux. In practice, the atomic beam flux is chosen to an appropriate finite value, taking into account the lifetime of the cesium clock. When the atomic beam flux is constant, the frequency stability of optically pumped cesium atomic clocks can be greatly improved by adopting a laser with narrow linewidth, low frequency noise, and high power stability, which can effectively improve the signal-to-noise ratio (SNR) of Ramsey fringes of the optically pumped cesium beam clock[17]. At the same time, miniaturization and transportability also need to be considered. The volume of the cesium beam clock is mainly limited by the cesium beam tube, but the shortening of the free drift region will reduce the Ramsey evolution time, which will widen the linewidth of the Ramsey fringes to a certain extent and lead to the the deterioration in frequency stability of cesium beam clock. Thus, the optimization of SNR can make the cesium beam clock still have excellent performance. Therefore, lasers with narrow linewidth and low frequency noise are essential.
@@
In order to optimize the frequency stability of optically pumped cesium beam clocks, we use frequency-stabilized Faraday laser [18, 19, 20, 21, 22, 23] as the pumping and detection laser sources. Compared with traditional external-cavity semiconductor lasers (ECDLs), such as lasers frequency selected by interferences and gratings, the Faraday laser has the advantages of immunity to temperature and current variations of the laser diode, and the frequency of Faraday laser can be automatically corresponding to atomic transition lines as soon as turning on, since Faraday laser is frequency selected by atomic filter[18, 19, 20]. Moreover, the modulation transfer spectroscopy (MTS) technique[24, 25, 26] is applied to lock the Faraday laser. Compared with other frequency stabilization methods, such as the saturation absorption spectroscopy (SAS)[27, 28, 29], polarization spectroscopy (PS)[30, 31], dichroic atomic vapor laser lock (DAVLL)[32, 33], MTS technique effectively suppress low-frequency noise, and it has the characteristics of high SNR and background-free[24].
@@
In this work, we achieve a high-performance Faraday laser pumped cesium beam clock, which applies an 852 nm Faraday laser frequency-stabilized by MTS technique as pumping and detection lasers. We measure the performance of the frequency-stabilized Faraday laser using the optical heterodyne method, and the result shows that the fractional frequency stability of each system is at 1 s, The most probable linewidth of the heterodyne beating signal is 12.7 kHz when Faraday lasers are free-running, which means that of the single system is 9.0 kHz. The most probable linewidth is 3.5 kHz with Faraday lasers being frequency locked by MTS, thus that of the single system is 2.5 kHz. We use such Faraday laser with narrow linewidth and low frequency noise as the pumping and the detection lasers to build the compact optically pumped cesium beam atomic clock, leading to the SNR of the Ramsey fringe being optimized. Thus, the fractional frequency stability of the Faraday laser pumped cesium beam clock can achieve at 1 s and at 100 s, dropping to at 10000 s when the cesium oven temperature is 110°C.
@@ II Experiment
Figure 1 shows the overall structure of the experiment. The Faraday laser is an integrated system with MTS technique for frequency stabilization, and all components are integrated onto a single aluminum base. The Faraday laser is beam-split and frequency-shifted to serve as the pumping and detection lasers for the cesium beam clock.
@@
The structure of the Faraday laser is depicted in Fig. 1. We employ an anti-reflection coated laser diode (ARLD) as the gain medium. The anti-reflection coating superbly suppresses the impact of the internal cavity modes, which are formed between the front and rear surfaces of the ARLD, on the laser modes. Two PBSs, a cesium vapor cell, and permanent magnets compose the Faraday anomalous dispersion optical filter (FADOF) which is for frequency selection. The vapor cell is filled with cesium atoms and 10 torr argon buffer gas, with a diameter of 1.5 cm and a length of 3 cm. The cesium vapor cell is surrounded with a twisted copper heating wire driven by a heater temperature controller and enclosed by thermal insulation materials. The temperature of the cesium vapor cell is controlled at 66°C with a temperature fluctuation of 0.1°C. The permanent magnets are placed around the cesium vapor cell to form a 900 G longitudinal magnetic field. Under the action of the magnetic field, the energy levels of the cesium atoms in the vapor cell undergo Zeeman splitting, which gives rise to the Faraday effect. The polarization direction of the linearly polarized light passing through the first PBS rotates as it propagates through the cesium vapor cell, and ultimately only the light whose polarization direction has rotated by can pass through the second PBS. Thus, only a very narrow frequency range of light corresponding to the spectral lines of the atoms can pass through the FADOF. Then the light is reflected back to the laser diode by a high-reflectivity mirror to form the external cavity feedback. Additionally, a piezoelectric transducer (PZT) is affixed to the high-reflectivity mirror, enabling the laser frequency to be tuned by adjusting the external cavity length. Due to the use of ARLD and the extremely narrow transmission spectrum of the FADOF, Faraday lasers are immune to the changes in current and temperature, a characteristic that distinguishes them from other types of ECDLs.
@@
In the MTS system, the frequency of the Faraday laser is locked to the cycling transition of by the MTS technique, as shown in Fig. 1. The Faraday laser beam is divided into pump laser and probe laser for MTS locking by a half wave plate (HWP) and a PBS. The pump laser is modulated by an electro-optic modulator (EOM) with a modulation frequency of 4.6 MHz. Then the modulation of the pump laser is transferred to the probe laser, which is received by the photoelectric detector (PD). The electrical signal of the PD is first amplified by an amplifier and then demodulated by a mixer with a demodulation frequency of 4.6 MHz to derive the MTS signal as shown in Fig. 2. The MTS error signal is processed by a proportional integral derivative (PID) circuit (Vescent D2-125) to form a feedback signal, which is sent to the laser controller and the laser diode respectively for frequency stabilization. The former signal is processed by the laser controller and then fed back to the current and piezoelectric ceramic (PZT) of the Faraday laser.
@@
As shown in Fig. 1, we implement two identical 852 nm frequency-stabilized Faraday laser systems to realize the optical heterodyne measurement. When both Faraday lasers are free-running, the signal derived by PD is sent into the spectrum analyzer, then we can obtain the fitted full width at half maximum (FWHM) of the heterodyne beating signal, as shown in Fig. 3(a). 30 sets of data are measured to get the most probable linewidth of the free-running Faraday laser which is statistically calculated to be 12.7 kHz, so that the most available linewidth for a single laser is 9.0 kHz, as shown in Fig. 3(b). When both laser systems are locked, the linewidth of the 852 nm MTS systems is obtained by the same method. The corresponding heterodyne beating signal obtained is shown in Fig. 3(c). The most probable linewidth between the two frequency-stabilized Faraday lasers is 3.5 kHz, so the most allowable linewidth of a single laser is 2.5 kHz, as shown in Fig. 3(d). The MTS technique greatly reduces the noise of the free-running Faraday laser. The narrow linewidth characteristic of the frequency-stabilized Faraday laser can significantly expand its application range.
@@
At the same time, the heterodyne beating signal from the PD is sent into a frequency counter (Keysight 53230A) to measure the frequency stability of the Faraday laser. The frequency stability of the free-running Faraday laser measured by the heterodyne method is at 1 s, as shown in Fig. 4. With both laser systems locked, the corresponding calculated frequency stability is shown in Fig. 4. The frequency stability of the heterodyne beating signal is at 1 s, so the frequency stability of the single frequency-stabilized Faraday laser is at 1 s. In the MTS system, we use a double-layer vapor cell as the absolute frequency reference. Changes in the temperature of the atomic vapor cell bring about a drift in the frequency of the cesium transition lines, whereas the space between the two quartz layers of the double-layer vapor cell is pumped into a vacuum, and the double-layer vapor cell is therefore well insulated from the effect of ambient temperature fluctuations on the frequency of the cesium transition lines.
@@
In the compact optically pumped cesium beam atomic clock, the pumping and the detection lasers are derived from the same Faraday laser frequency-stabilized on the cesium cycling transition. The output of the Faraday laser is split into two beams. One beam is used directly as the detection laser, while the other beam is frequency-shifted by 251.1 MHz using an acousto-optic modulator (AOM) and subsequently used as the pumping laser, which has a frequency corresponding to the transition line. The corresponding atomic transitions are schematically shown in Fig 1. Both the pumping and the detection lasers pass through the beam expanders before being injected into the cesium beam tube, enlarging the interaction area between the laser beams and the cesium beam by a factor of . In the area where the cesium beam interacts with the pumping laser, a reflector is mounted on the back of the cesium beam tube to reflect the pumping laser to interact with the cesium beam again to improve the pumping efficiency.
@@ III Results
A microwave signal synthesized by a 10 MHz voltage-controlled crystal oscillator (VCXO) is acted on the microwave cavity and tuned with a center frequency of 9.192631770 GHz. The cesium oven temperature is set at 110 °C. The linewidth and the amplitude of the Ramsey fringe is 551.4 Hz and 461.85 mV, respectively, as shown in Fig. 5. The cesium beam clock operates at a modulation frequency of 137 Hz, and the noise of the Ramsey fringe signal is analyzed by a fast Fourier transform spectrum analyzer (Stanford Research Systems, SR770). The noise at 137 Hz is . Thus, the SNR in 1 Hz bandwidth at 137 Hz is given by , where represents the amplitude of the Ramsey fringe and represents the noise of the Ramsey fringe at 137 Hz. The Ramsey fringe signal is modulated and demodulated to produce an error signal which is used to stabilize the output frequency of the VCXO by the servo circuits.
After locking the 10 MHz VCXO of the cesium beam clock, we compare the 10 MHz output of the VCXO with a hydrogen maser (VREMYA-CH, VCH-1003M, option L) by a high-performance phase noise and Allan deviation analyzer (Microsemi, 5120A). The frequency stability of the Faraday laser pumped cesium beam clock is two times better than that of the DFB laser-based cesium beam clock[34], as shown in Fig. 6. In addition, the frequency stability of our cesium beam clock is also significantly better than that of the cesium beam clock based on the interferometer ECDL[13], which is at 100 s. The three cesium beam clocks are different only in the categories of the lasers. That is, they use the same cesium beam tube and electronic system, and their cesium oven temperatures are all set at 100°C. Thus, the comparison results highlight the superiority of the Faraday laser, which has narrow linewidth and high frequency stability.
We try different cesium oven temperatures, which would directly affect the SNR of the Ramsey fringe. As shown in Fig. 7, when the cesium oven temperature is increased to 110°C, the frequency stability of the cesium beam clock reaches , decreasing to at 10000 s. The frequency stability of our clock is better than that of any reported compact cesium beam atomic clocks to our knowledge, which is attributed to the narrow linewidth and the low frequency noise of the Faraday laser system resulting in a high SNR of the Ramsey fringe. Meanwhile, the frequency stability of our clock is at 1 s , which firstly reaches the magnitude. When the cesium oven temperature is 130°C, the SNR of the Ramsey fringe in 1 Hz bandwidth at 137 Hz will reach , which can result in better frequency stability. Taking into account the lifetime of the cesium beam tube, the oven temperature is set at 110°C. The increase in signal intensity brought about by the increase in oven temperature does not necessarily lead directly to a higher SNR, which requires the support of a laser with low frequency noise[35].